OpenStudy (anonymous):

Alex designed the model of a cylindrical water tank to be installed in a park. The model has a radius of 5 cm and a height of 14 cm. The dimensions of the actual water tanks that will be installed in the park are 8 times larger than the model. The total surface area of the actual water tanks, rounded to the nearest whole number, is _________ cm2. use 22/ 7 for pi .

OpenStudy (anonymous):

@monokerous

OpenStudy (anonymous):

@AccessDenied

OpenStudy (anonymous):

formula for area of cylinder:- 2pi.r^2+2pi.rh surface area for model cylinder:- 2*22/7*5*5+ 2*22/7*25*14=157.14+2200 =2357.14cm^2 surface area for larger model:- 2357.14*8 =18857.12cm^2

OpenStudy (anonymous):

=18857

OpenStudy (anonymous):

Okay , i think i get it now . Thanks a bunch